OpenStudy (anonymous):

I need help solving this problem. √-225 + √-49

OpenStudy (anonymous):

15i +7i=22i

OpenStudy (anonymous):

\[\sqrt{-225} = \sqrt{-1} * \sqrt{225}\] = 25i Look at it that way

OpenStudy (anonymous):

as roots can't be defined for negative numbers, they are "imaginary" and the imaginary number is "-i"

OpenStudy (anonymous):

*15i

OpenStudy (anonymous):

yeah

OpenStudy (radar):

dont forget to ads the 7i you got from sq root of -49! to get the total of 22i

OpenStudy (anonymous):

lol...the answer is 15i+7i

OpenStudy (anonymous):

Or \[\sqrt{-484}\]

OpenStudy (radar):

Krb686. Please show how \[\sqrt{-225}+\sqrt{-49}=\sqrt{-484}\]??? Enquiring minds want to know!

OpenStudy (anonymous):

Because 15i + 7i = 22i \[i = \sqrt{-1}\] \[22*\sqrt{-1}\] \[\sqrt{22^{2}*-1}\]

OpenStudy (radar):

Thanks, enquiring mind now happy lol.

OpenStudy (anonymous):

:D